From GlueXWiki
Jump to: navigation, search

January 26, 2012 FDC meeting


  1. Production update Construction Tracking (Dave)
  2. First package testing status (Lubomir, Beni)
  3. Engineering update (Bill)
  4. Electronics update (Chris)
  5. Other


Participants: Bill, Dave, Chris, Nick, Mark, Simon, Beni, and Lubomir.


- Dave: wire board #20 (first cell of the 3rd package) is ready for wire deadening; waiting for the microscope results from Olga to make sure the new timing/currents for electroplating is OK. Wire board #21 ready for tension measurements, wire board #22 ready for stringing, wire board #23 was laminated. Cathodes: the two type-3 two tyoe-1 and one type-2 ready.

- Casey is modifying the frame to install the second package on the rotational brackets (taken from the first package, see discussions later). After the second package is moved the tooling needed for the third package will be installed in the testing room. It includes 14 hubs and three metal pins to be taken from package #2. Package number #2 will remain with three metal pins and plastic rods only.

First package testing status

- Lubomir: First package moved to EEL 126 on Monday; some problems discussed below.

- Gas problem: initially it was bubbling at >220ccpm (compared to 40-50ccpm at Blue Crab), but there was no obvious gas leakage except some (1x10^-4 on the leakage detector scale) at the southern side where the input gas connectors were. Today the input and output connectors were swapped and the problem almost disappeared, which indicates the leakage was somewhere at gas connectors or close that to in the chamber. Has to be found and fixed later!

- HV module problem: the CAEN HV supply module (A1535) has current resolution of 500nA, to be compared to the dark current of ~10-100nA per HV sector. That makes you blind when working/conditioning the chambers since you don't see if the cables are connected or if the current oscillates. At least there are no problems with the safety of the chamber: you can apply the minimal threshold of 500nA and even if it trips at 1-2uA it is safe. There are CAEN modules (A1550) with the same connector configuration with 100nA resolution and even special for chambers with 5nA resolution (A2932). It's Lubomir's and Beni's fault, we didn't look carefully at the PR specifications in 2010, but same modules are used by Hall B in much bigger quantities!? Lubomir will discuss with Fernando and Eugene about possible solutions.

- Other HV problems so far: two negative and two positive HV channels are tripping even if 1 Volt is applied. Could be the module, the cable, or connector problem; easy to identify.

- Other than that the package is now bubbling with 60/40 gas mixture and the HV so far is 1900/-500V on all the channels except the faulty ones and few more are at 1800. We will continue with the conditioning, and at the same time will start looking at the signals.

- Beni: the MKS gas system is not reliable and turns off several times a day which makes testing almost impossible. Beni wants to start building the new system that will be used in the Hall. For the beginning his plan is to use the new controllers and asked Bill to design mixing volume with some tubing, that is the minimal equipment needed for the tests. Still waiting for three additional F1TDCs needed to instrument all the wire planes.


- The way the strong-back is attached to the chamber is not the best: we want the holding brackets to be attached directly to the strong-back, not to the chamber that holds the strong-back now. This also may be related to the gas problem. Bill proposed to install beams bellow the strong-back and put the package sitting on them while testing in 126. At the same time he will look for a way to attach the brackets to the strong-back that will be needed when working on the HV cabling and transporting/carrying the package.

- Cooling many-folds are designed, pumps have arrived; Bill plans to start testing parts of the cooling system in 126, including installation on the first package. The additional cooling brackets for the pre-amps arrived; Bill finished also some assemby drawings. The tooling for LV connector installation was tested after the meeting: can't be used at some places and requires modifications.


- Chris: three LV cable and three jumpers (enough for all wire planes) are ready. Difficult to install LV connectors on these wires. Maybe we will not need in the near future the cathode LV cables. It turns out that for grounding the strips, we don't need LV on the cathode cards. That was measured by Fernando and the signals from 55Fe source with cathode and grounding cards were compared by Lubomir: almost no difference. Though, the noise looked bigger with the cathode cards and this has to be investigated on the first package.